# The Memorization Problem: Can We Trust LLMs’ Economic Forecasts?

### v1

@@ 4.2 Market Index Recall
We next evaluate GPT-4o’s memorization of market index data by testing its ability to recall daily and monthly values for the S&P 500, Dow Jones Industrial Average (DJIA), and Nasdaq Composite, using data from December 1989 to March 2025. For numerical recall tests, we requested exact closing values at daily frequency, both without context and with the previous two days’ levels provided, as well as monthly returns. Additionally, we assessed directional changes (up or down) and relative performance between index pairs at monthly frequency. Performance metrics include Mean Percent Error, Mean Absolute Percent Error, and Directional Accuracy Change (proportion of predictions correctly identifying the direction of change relative to the previous period) for numerical predictions and accuracy for directional and relative performance tasks, all compared against actual values from Yahoo Finance. Results are reported in Table 2, with pre-cutoff (before October 2023) and post-cutoff (after October 2023) periods distinguished to isolate memorization effects.
Table 2: Evaluation Metrics for Market Indices
This table reports a set of evaluation metrics assessing the LLM’s ability to recall market index levels and their changes over time. These tests are done at the daily or monthly frequency. We ask the LLM to recall the closing value of the index each trading day. Panel A provides metrics for predictions of Daily Levels and Daily Levels with context (where the previous two days’ index levels are provided). We ask the LLM to provide monthly returns for these indices as well. Metrics include Mean Percent Error (MPE), Mean Absolute Percent Error (MAPE), and Directional Accuracy Change, all reported in percentage points (0.10 means 0.10%). MPE is calculated by averaging the percent error $(EstimatedLevel-ActualLevel)/ActualLevel$. MAPE takes the average absolute value of the percent errors. Directional Accuracy Change measures the proportion of predictions correctly identifying the direction of change (up or down) relative to the previous day. Confidence Calibration reports the correlation between the LLM’s confidence level (on a scale from 0 to 100) and mean absolute percent error. Panel B presents accuracy metrics related to predicting Directional Changes and Relative Performance between indices. Directional Changes asks the LLM directly for an up or down answer for each month. Relative Performance asks the LLM to answer which index of the index pair performed better during the month. Accuracy reports the proportion of predictions correctly identifying either the direction of change or relative performance in percentage points. Confidence Calibration in this panel reflects the correlation between the LLM’s confidence and the MAPE. Results are separately provided for the S&P 500 (SP500), Dow Jones Industrial Average (DJIA), and Nasdaq Composite indices.
@@
MPE (%) MAPE (%) Directional Accuracy Change (%) Confidence Calibration Num Obs Start Date End Date Refusals
Daily Levels: Pre-cutoff
SP500 0.12 0.61 80.68 -0.14 8489 12/29/1989 09/29/2023 0
DJIA 0.10 0.47 82.23 -0.35 7982 01/02/1992 09/29/2023 0
Nasdaq Composite 0.18 1.80 69.44 -0.12 8489 12/29/1989 09/29/2023 0
Daily Levels: Post-cutoff
SP500 -16.59 16.70 45.97 -0.13 249 10/02/2023 03/05/2025 64
DJIA -12.97 13.03 49.81 -0.15 270 10/02/2023 03/04/2025 87
Nasdaq Composite -19.66 19.75 44.10 -0.40 230 10/02/2023 03/05/2025 62
Daily Levels: Pre-cutoff with context
SP500 0.13 0.50 80.84 -0.16 8497 12/29/1989 09/29/2023 0
DJIA 0.05 0.43 81.60 -0.17 7983 01/06/1992 09/29/2023 510
Nasdaq Composite 0.00 1.06 68.77 -0.18 8492 12/29/1989 09/29/2023 2
Daily Levels: Post-cutoff with context
SP500 0.02 0.63 57.09 0.03 290 10/03/2023 03/07/2025 36
DJIA -12.97 13.03 49.81 -0.15 270 10/02/2023 03/04/2025 87
Nasdaq Composite -19.66 19.75 44.10 -0.40 230 10/02/2023 03/05/2025 62
Monthly Returns: Pre-cutoff
SP500 -0.70 3.35 85.26 0.27 407 12/01/1989 10/01/2023 30
DJIA -0.70 3.25 81.10 0.29 381 02/01/1992 10/01/2023 25
Nasdaq Composite -1.01 4.81 79.36 0.21 407 12/01/1989 10/01/2023 44
@@
For pre-cutoff daily exact numerical levels, GPT-4o exhibits strong recall, with Mean Absolute Percent Errors of 0.61% for S&P 500, 0.47% for DJIA, and 1.80% for Nasdaq Composite, and Directional Accuracy Change ranging from 69.44% (Nasdaq) to 82.23% (DJIA). Providing context improves slightly accuracy for S&P 500 (0.58%) and Nasdaq (1.06%). Prompting directly for returns, which we test monthly, yields higher Directional Accuracy (79.36%–85.26%), reflecting robust memorization of directional trends. Other tests further confirm memorization: prompting directly for directional performance (”up” or ”down”) exceeds 79% accuracy across indices, and relative performance accuracy ranges from 82.86% (S&P 500 vs. Nasdaq) to 87.10% (S&P 500 vs. DJIA).
In contrast, post-cutoff performance collapses, with Mean Absolute Percent Errors ballooning to 13.03%–19.75% for exact levels and Directional Accuracy Change dropping to near-random levels (44.10%–49.81%), indicating no data leakage beyond the training cutoff.
The sharp pre-cutoff accuracy, particularly for exact levels and relative performance, highlights GPT-4o’s extensive memorization of historical index data, posing challenges for forecasting studies. The model’s ability to recall precise closing values and correctly identify directional trends within its training period suggests that any apparent predictive success may reflect memorized data rather than analytical capability. The negligible improvement from context and the complete performance drop post-cutoff reinforce that these results stem from training data exposure. These findings caution against using LLMs for historical market analysis without ensuring data is outside their training scope, as their outputs risk being artifacts of memorization rather than genuine economic foresight.
@@ 4.3 Headline Date Identification
For headline date identification, we present GPT-4o with sets of Wall Street Journal front page headlines (approximately 9 headlines per day) from our dataset of 90,123 headlines spanning December 1989 to February 2025 without revealing their publication dates. We asked the model to identify when these headlines appeared and, in a separate test variant, predict the S&P 500 level on the following trading day. Performance was evaluated using multiple accuracy metrics: year accuracy, month-and-year accuracy, exact date accuracy, mean absolute days difference, and confidence calibration. By comparing results between pre-training headlines (where memorization could occur) and post-training headlines (where memorization is impossible), we can clearly distinguish between the model’s inferential abilities and its capacity to recall memorized chronological information.
Table 3: Evaluation Metrics for Headlines
@@
Pre-training Period -0.77 9.52 98.45 90.38 47.03 -0.10
Post-training Period 413.46 414.54 28.81 20.71 7.86 -0.12
@@
We present the results in Table 3. GPT-4o demonstrates remarkable memorization of headline chronology within its training period. For pre-cutoff headlines, it achieves 98.45% accuracy in determining the correct year and 90.38% in identifying the correct month and year. Even for exact date identification, the model achieves 47.03% accuracy—significantly above chance levels. When incorrect, the model’s estimates remain close to the actual date, with a mean absolute difference of 9.52 days.
In stark contrast, for headlines published after the model’s training cutoff date, performance deteriorates dramatically across all metrics. Year accuracy drops to 28.81%, month-and-year accuracy falls to 20.71%, and exact date accuracy declines to just 7.86%. The mean absolute difference increases to 414.54 days, indicating essentially random guessing.
We observed a similar pattern when we extended our test to ask the model to provide both the headline date and the corresponding S&P 500 level on the next trading day. For the pre-training period, the model achieved high temporal accuracy while maintaining near-perfect recall of index values (mean absolute percent error of just 0.01%). For post-training headlines, both date identification and index level predictions became significantly less accurate.
The sharp performance discontinuity at the training cutoff date provides compelling evidence that the model’s apparent ”knowledge” of financial chronology stems primarily from memorization rather than inference or reasoning. This finding raises significant concerns about using LLMs to analyze historical relationships between news events and market movements within their training period, as their responses may reflect memorized associations rather than genuine analytical insights.
@@ 4.4 Individual Stock Price Recall
To further investigate GPT-4o’s memorization capabilities, we tested its ability to recall end-of-month closing prices for the Magnificent 7 stocks (META, GOOGL, AMZN, TSLA, NVDA, MSFT, AAPL) from January 1989 to September 2023, all within the model’s training cutoff of October 2023. We queried prices both without context and with the previous two months’ closing prices provided, using data from the Center for Research in Security Prices (CRSP). Performance was evaluated using Mean Percent Error, Mean Absolute Percent Error, and Directional Accuracy (correctly identifying the direction of change relative to the previous month), with results compared against actual closing prices. Table 4 reports these metrics, complementing our earlier findings on market indices by examining individual security-level memorization. We also plot the actual vs estimated values in Figures 4 and 5.
@@
This table reports a set of evaluation metrics for the Magnificent 7 stocks which includes META, GOOGL, AMZN, TSLA, NVDA, MSFT, and AAPL. We ask the LLM to recall closing prices at the end of each month. Mean Percent Error (MPE), Mean Absolute Percent Error (MAPE), and Directional Accuracy Change are reported in percentage points (0.18 means 0.18%). MPE is calculated by taking the average of the percent error $(PredictedPrice-ActualPrice)/ActualPrice$. MAPE is calculated by taking the average of the absolute value of the percent error. Directional Accuracy Change is the proportion of predictions that went in the correct direction (up or down) with respect to the previous month. Confidence Calibration is the correlation between the LLM’s confidence level (on a scale of 0 to 100) and the MAPE. Num Obs is the number of observations used in the evaluation, Start Date and End Date indicate the period over which the metrics were computed. Refusals are the number of instances in which the model withheld a prediction by either answering ”null” or 0. Results are provided for a prompt that contains an empty context in panel A and a prompt that provides the previous two month’s closing prices as context in panel B.
@@
MPE (%) MAPE (%) Directional Accuracy Change (%) Confidence Calibration Num Obs Start Date End Date Refusals
META 0.18 0.37 99.26 -0.08 137 05/31/2012 09/29/2023 0
GOOGL -1.41 1.79 93.42 -0.19 229 08/31/2004 09/29/2023 1
AMZN -5.87 7.98 91.77 -0.12 317 05/30/1997 09/29/2023 0
TSLA -9.21 9.99 92.45 -0.13 160 06/30/2010 09/29/2023 0
NVDA -20.60 23.92 77.05 -0.53 293 01/29/1999 09/29/2023 4
MSFT -26.09 26.98 76.62 -0.66 403 01/31/1989 09/29/2023 14
AAPL -36.94 38.21 72.68 -0.57 411 01/31/1989 09/29/2023 6
@@
MPE (%) MAPE (%) Directional Accuracy Change (%) Confidence Calibration Num Obs Start Date End Date Refusals
META -0.18 0.40 98.52 -0.10 136 05/31/2012 09/29/2023 1
GOOGL -0.27 0.84 95.52 0.10 224 08/31/2004 09/29/2023 6
AMZN -0.42 3.12 93.35 0.02 317 05/30/1997 09/29/2023 0
TSLA -0.87 2.34 94.87 -0.17 157 06/30/2010 07/31/2023 3
NVDA 1.01 8.80 80.68 -0.17 296 01/29/1999 08/31/2023 1
MSFT 1.24 4.57 84.71 -0.21 400 01/31/1990 07/31/2023 5
AAPL -1.37 5.89 83.91 -0.25 405 01/31/1990 09/29/2023 0
@@
The results reveal varying recall accuracy across stocks, with notable improvements when context is provided. Without context (Panel A), GPT-4o performs best for newer stocks like META, with a Mean Absolute Percent Error of 0.37% and Directional Accuracy Change of 99.26%, but struggles with older stocks like AAPL (38.21% error, 72.68% accuracy) and MSFT (26.98% error, 76.62% accuracy). Errors are also high for NVDA (23.92%) and TSLA (9.99%), suggesting selective memorization tied to stock age or data prominence. With context (Panel B), accuracy improves significantly: Mean Absolute Percent Errors drop to 0.40% for META, 0.84% for GOOGL, and 5.89% for AAPL, with Directional Accuracy Change rising to 98.52%, 95.52%, and 83.91
Figure 4: Recall of closing prices for META, GOOGL, AMZN, and TSLA.
This figure shows the LLM’s estimated closing prices for META, GOOGL, AMZN, and TSLA compared to the actual values. Panels A, C, E, and G graph the actual values against the estimated values. Panels B, D, F, and H show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%).
Figure 5: Recall of closing prices for NVDA, AAPL, and MSFT.
This figure shows the LLM’s estimated closing prices for NVDA, AAPL, and MSFT compared to the actual values. Panels A, C, and E graph the actual values against the estimated values. Panels B, D, and F show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%).
Figure 6: Recall of closing prices for NVDA, AAPL, and MSFT with context.
This figure shows the LLM’s estimated closing prices for NVDA, AAPL, and MSFT compared to the actual values. We give the LLM two previous end of the month closing prices given as context. Panels A, C, and E graph the actual values against the estimated values. Panels B, D, and F show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%).
These findings extend our market index results, highlighting that GPT-4o’s memorization is not uniform across securities and is sensitive to contextual cues and stock-specific factors. The high accuracy for META and GOOGL, especially with context, parallels the model’s strong recall of recent macroeconomic levels and index values, suggesting robust memorization of prominent, frequently referenced data. Conversely, larger errors for older stocks like AAPL and MSFT, even with context, align with the weaker recall of long-horizon macroeconomic levels, pointing to the potential dilution of older data in the training corpus. This selective memorization reinforces the challenge for financial research: apparent forecasting success for individual stocks within the training period may reflect memorized prices rather than predictive ability, necessitating evaluations with post-cutoff data to ensure methodological rigor.
@@
Moreover, Figure 6 shows why any memorization performance only provides a lower bound. The errors substantially decrease when we give ChatGPT the prices for the previous two months. This situation is relevant as when forecasting, researchers typically provide contextual information.
4.5 Portfolio Stock Price Recall by Size

### v2

@@ 5.2 Market Index Recall
We next evaluate GPT-4o’s memorization of market index data by testing its ability to recall daily and monthly values for the S&P 500, Dow Jones Industrial Average (DJIA), and Nasdaq Composite, using data from January 1990 to February 2025. For numerical recall tests, we requested exact closing values at daily frequency, both without context and with the previous two days’ levels provided, as well as monthly returns. Additionally, we assessed directional changes (up or down) and relative performance between index pairs at the monthly frequency. Performance metrics include Mean Percent Error, Mean Absolute Percent Error, and Directional Accuracy (proportion of predictions correctly identifying the direction of change relative to the previous period) for numerical predictions and accuracy for directional and relative performance tasks, all compared against actual values.
Table 2: Evaluation Metrics for Stock Market Indices
This table reports a set of evaluation metrics assessing the LLM’s ability to recall market index levels and their changes over time. These tests are done at the daily or monthly frequency. We ask the LLM to recall the closing value of the index each trading day. Panel A provides metrics for predictions of Daily Levels and Daily Levels with context (where the previous two days’ index levels are provided). We ask the LLM to provide monthly returns for these indices as well. Metrics include Mean Percent Error (MPE), Mean Absolute Percent Error (MAPE), and Directional Accuracy, all reported in percentage points (0.10 means 0.10%). MPE is calculated by averaging the percent error (EstimatedLevelActualLevel)/ActualLevel(EstimatedLevel-ActualLevel)/ActualLevel. MAPE takes the average absolute value of the percent errors. Directional Accuracy measures the proportion of predictions of the market index levels correctly following the direction of change (up or down) relative to the previous day. Confidence Calibration reports the correlation between the LLM’s confidence level (on a scale from 0 to 100) and mean absolute percent error. Panel B presents accuracy metrics related to predicting Directional Changes and Relative Performance between indices. Directional Changes asks the LLM directly for an up or down answer for each month. Relative Performance asks the LLM to answer which index of the index pair performed better during the month. Accuracy reports the proportion of predictions correctly identifying either the direction of change or relative performance in percentage points. Confidence Calibration in this panel reflects the correlation between the LLM’s confidence and the MAPE. Results are separately provided for the S&P 500 (SP500), Dow Jones Industrial Average (DJIA), and Nasdaq Composite indices.
@@
MPE (%) MAPE (%) Directional Accuracy (%) Confidence Calibration Num Obs Refusals
Daily Levels: Pre-cutoff, 01/02/1990 to 09/29/2023
SP500 0.12 0.61 80.58 -0.14 8,488 0
DJIA 0.11 0.53 80.66 -0.36 8,488 0
Nasdaq Composite 0.18 1.80 69.38 -0.12 8,488 0
Daily Levels: Post-cutoff, 10/02/2023 to 02/28/2025
SP500 -16.78 16.87 45.70 -0.10 292 62
DJIA -12.96 13.01 49.26 -0.14 271 83
Nasdaq Composite -20.40 20.48 44.03 -0.25 294 60
Daily Levels: Pre-cutoff with context, 01/02/1990 to 09/29/2023
SP500 0.13 0.50 80.78 -0.16 8,488 0
DJIA 0.06 0.46 80.06 -0.17 8,488 0
Nasdaq Composite 0.00 1.06 68.62 -0.18 8,488 0
Daily Levels: Post-cutoff with context, 10/02/2023 to 02/28/2025
SP500 0.00 0.64 54.40 -0.07 319 35
DJIA -0.04 0.54 53.17 0.02 332 22
Nasdaq Composite -0.03 0.92 51.85 -0.09 298 56
Monthly Returns: Pre-cutoff, 01/1990 to 09/2023
SP500 -0.70 3.36 85.19 0.27 405 0
DJIA -0.70 3.28 80.49 0.26 405 0
Nasdaq Composite -1.03 4.83 79.51 0.21 405 0
Monthly Returns: Post-cutoff, 10/2023 to 02/2025
SP500 -1.16 3.20 41.67 -0.28 12 5
DJIA -0.80 3.18 30.77 -0.06 13 4
Nasdaq Composite -2.46 4.07 41.67 -0.39 12 5
@@
Results are reported in Table 2, distinguishing pre-cutoff (before October 2023) and post-cutoff (after October 2023) periods to isolate memorization effects. For pre-cutoff daily exact numerical levels, GPT-4o exhibits strong recall, with Mean Absolute Percent Errors of 0.61% for S&P 500, 0.53% for DJIA, and 1.80% for Nasdaq Composite, and Directional Accuracy ranging from 69.4% (Nasdaq) to 80.7% (DJIA). Providing context improves slightly accuracy for S&P 500 (0.50%) and Nasdaq (1.06%). Prompting directly for returns, which we test monthly, yields higher Directional Accuracy (79.5%–85.2%), reflecting robust memorization of directional trends. Other tests further confirm memorization: prompting directly for directional performance (“up” or “own”) exceeds 79% accuracy across indices, and relative performance accuracy ranges from 82.9% (S&P 500 vs. Nasdaq) to 87.1% (S&P 500 vs. DJIA).
In contrast, post-cutoff performance collapses. For instance, without context, the Mean Absolute Percent Errors balloon to 13.01%–20.48% for exact levels and Directional Accuracy dropping to near-random levels (44.0%–49.3%). We find similar evidence for recall with context in terms of Directional Accuracy which ranges from 51.9% to 54.4% in the post-cutoff period, much lower than the Directional Accuracy pre-cutoff which ranges from 68.6% to 80.8%. These findings do not suggest data leakage beyond the training cutoff date.
The sharp pre-cutoff accuracy, particularly for exact levels and relative performance, highlights GPT-4o’s extensive memorization of historical index data, posing challenges for forecasting studies in economics and finance. The model’s ability to recall precise closing values and correctly identify directional trends within its training period suggests that any apparent predictive success may reflect memorized data rather than analytical capability. The negligible improvement from context and the complete performance drop post-cutoff reinforce that these results stem from training data exposure. These findings caution against using LLMs for historical market analysis without ensuring data is outside their training scope, as their outputs risk being artifacts of memorization rather than genuine economic foresight.
@@ 5.3 Headline Date Identification
In this section, we assess LLMs’ ability to extract date information from front-page news headlines and, in turn, predict the aggregate stock market index. To do so, we present GPT-4o with sets of The Wall Street Journal front-page headlines (approximately 9 headlines per day) from our dataset of 90,123 headlines spanning January 1990 to February 2025 without revealing their publication dates. We asked the model to identify when these headlines appeared and, in a separate test variant, predict the S&P 500 level on the next trading day. Performance is evaluated using multiple accuracy metrics: year accuracy, month-and-year accuracy, exact date accuracy, mean absolute days difference, and confidence calibration. By comparing results between pre-cutoff headlines (where memorization could occur) and post-cutoff headlines (where memorization is impossible), we can clearly distinguish between the model’s inferential abilities and its capacity to recall memorized chronological information.
Table 3: Evaluation Metrics for News Headlines
@@
Pre-training Period -0.77 9.52 98.45 90.38 47.03 -0.10 - -
Post-training Period 413.46 414.54 28.81 20.71 7.86 -0.12 - -
@@
We present the results in Table 3. GPT-4o demonstrates remarkable memorization of headline chronology within its training period. For pre-cutoff headlines, it achieves 98.5% accuracy in determining the correct year and 90.4% in identifying the correct month and year. Even for exact date identification, the model achieves 47.0% accuracy—significantly above chance levels. When incorrect, the model’s estimates remain close to the actual date, with a mean absolute difference of 9.5 days.
In stark contrast, for headlines published after the model’s training cutoff date, performance deteriorates dramatically across all metrics. Year accuracy drops to 28.8%, month-and-year accuracy falls to 20.7%, and exact date accuracy declines to just 7.9%. The mean absolute difference increases to 414.5 days, indicating essentially random guessing.
We observed a similar pattern when we extend our test to ask the model to provide both the headline date and the corresponding S&P 500 level on the next trading day. For the pre-training period, the model achieves high temporal accuracy while maintaining near-perfect recall of index values (mean absolute percent error of just 0.01%). For post-training headlines, both date identification and index level predictions became significantly less accurate.
The sharp performance discontinuity at the training cutoff date provides compelling evidence that the model’s apparent “knowledge” of financial chronology stems primarily from memorization rather than inference or reasoning. This finding raises significant concerns about using LLMs to analyze historical relationships between news events and market movements within their training period, as their responses may reflect memorized associations rather than genuine analytical insights.
@@ 5.4 Individual Stock Price Recall
To complement our earlier findings on market indices, we examine memorization at the individual-security level in this section. In particular, we test the model’s ability to recall end-of-month closing prices for prominent stocks such as Magnificent 7 and individual stocks in the cross section.
5.4.1 Magnificent 7 stocks
We begin our analysis with the Magnificent 7 stocks (META, GOOGL, AMZN, TSLA, NVDA, MSFT, AAPL) from January 1990 to September 2023, which precedes the model’s training cutoff of October 2023. We queried prices both without context and with the previous two months’ closing prices provided, using data from the Center for Research in Security Prices (CRSP). Performance was evaluated using Mean Percent Error, Mean Absolute Percent Error, and Directional Accuracy (correctly identifying the direction of change relative to the previous month), with results compared against actual closing prices. We report these metrics in Table 4 and plot the actual vs estimated values without context in Figures 4 and 5.
@@ 5.4.1 Magnificent 7 stocks
This table reports a set of evaluation metrics for the Magnificent 7 stocks which includes META, GOOGL, AMZN, TSLA, NVDA, MSFT, and AAPL. We ask the LLM to recall closing prices at the end of each month. Mean Percent Error (MPE), Mean Absolute Percent Error (MAPE), and Directional Accuracy are reported in percentage points (0.18 means 0.18%). MPE is calculated by taking the average of the percent error (PredictedPriceActualPrice)/ActualPrice(PredictedPrice-ActualPrice)/ActualPrice. MAPE is calculated by taking the average of the absolute value of the percent error. Directional Accuracy is the proportion of predictions that went in the correct direction (up or down) with respect to the previous month. Confidence Calibration is the correlation between the LLM’s confidence level (on a scale of 0 to 100) and the MAPE. Num Obs is the number of observations used in the evaluation, Start Date and End Date indicate the period over which the metrics were computed. Refusals are the number of instances in which the model withheld a prediction by either answering ”null” or 0. Results are provided for a prompt that contains an empty context in panel A and a prompt that provides the previous two month’s closing prices as context in panel B.
@@
MPE (%) MAPE (%) Directional Accuracy (%) Confidence Calibration Num Obs Start Date End Date Refusals
Pre-cutoff, 01/1990 to 09/2023
META 0.18 0.37 99.26 -0.08 137 05/2012 09/2023 0
GOOGL -1.41 1.79 93.42 -0.19 229 08/2004 09/2023 1
AMZN -5.87 7.98 91.77 -0.12 317 05/1997 09/2023 0
TSLA -9.21 9.99 92.45 -0.13 160 06/2010 09/2023 0
NVDA -20.60 23.92 77.05 -0.53 293 01/1999 09/2023 4
MSFT -25.72 26.62 76.75 -0.65 401 01/1990 09/2023 14
AAPL -35.14 36.44 72.61 -0.54 399 01/1990 09/2023 6
Post-cutoff, 10/2023 to 02/2025
META -29.59 29.60 50.00 0.48 5 10/2023 02/2025 12
GOOGL 138.06 170.09 55.56 -0.98 10 10/2023 02/2025 7
AMZN -22.06 22.06 42.86 0.48 8 10/2023 02/2025 7
TSLA -0.81 20.15 100.00 0.61 6 10/2023 02/2025 11
NVDA 436.43 436.43 100.00 -0.64 8 10/2023 02/2025 9
MSFT -19.25 19.25 45.45 -0.48 12 10/2023 02/2025 5
AAPL -15.46 15.46 30.00 0.36 11 10/2023 02/2025 6
@@
MPE (%) MAPE (%) Directional Accuracy (%) Confidence Calibration Num Obs Start Date End Date Refusals
Pre-cutoff, 01/1990 to 09/2023
META -0.18 0.40 98.52 -0.10 136 05/2012 09/2023 1
GOOGL -0.27 0.84 95.52 0.10 224 08/2004 09/2023 6
AMZN -0.42 3.12 93.35 0.02 317 05/1997 09/2023 0
TSLA -0.87 2.34 94.87 -0.17 157 06/2010 09/2023 3
NVDA 1.01 8.80 80.68 -0.17 296 01/1999 09/2023 1
MSFT 1.24 4.57 84.71 -0.21 400 01/1990 09/2023 5
AAPL -1.37 5.89 83.91 -0.25 405 01/1990 09/2023 0
Post-cutoff, 10/2023 to 02/2025
META -5.04 5.29 50.00 0.14 7 10/2023 02/2025 10
GOOGL 0.88 6.26 50.00 -0.36 9 10/2023 02/2025 8
AMZN -2.72 6.06 54.55 0.54 12 10/2023 02/2025 5
TSLA -5.67 10.75 60.00 0.02 11 10/2023 02/2025 6
NVDA 85.89 99.39 66.67 0.45 10 10/2023 02/2025 7
MSFT -0.80 2.75 50.00 -0.14 9 10/2023 02/2025 8
AAPL -2.90 4.12 55.56 -0.10 10 10/2023 02/2025 7
@@
The results reveal varying recall accuracy across stocks, with notable improvements when context is provided. Without context (Panel A), GPT-4o performs best on more recently listed stocks, such as META, with a Mean Absolute Percent Error of 0.37% and Directional Accuracy of 99.26%, but struggles with stocks with a longer history, such as AAPL (36.44% error, 72.61% accuracy) and MSFT (26.62% error, 76.75% accuracy). Errors are also high for NVDA (23.92%) and TSLA (9.99%), suggesting selective memorization tied to the length of the price series or data prominence. With context, accuracy improves significantly as shown in Panel B: Mean Absolute Percent Errors drop to 0.40% for META, 0.84% for GOOGL, and 5.89% for AAPL, with Directional Accuracy rising to 98.52%, 95.52%, and 83.91%, respectively. This context-driven enhancement mirrors the slight improvements seen for market indices, indicating that recent price cues help the model anchor its recall more precisely, particularly for stocks with longer price histories.
@@
We plot the actual vs estimated values with context for NVDA, AAPL, and MSFT in Figure 6. Compared to Figure 5, this figure shows why any memorization performance only provides a lower bound. The errors substantially decrease when we give ChatGPT the prices for the previous two months. This situation is relevant as when forecasting, researchers typically provide contextual information.
Figure 4: Recall of closing prices for AMZN, GOOGL, META, and TSLA without context.
This figure shows the LLM’s estimated closing prices for AMZN, GOOGL, META, and TSLA compared to the actual values. Panels A, C, E, and G graph the actual values against the estimated values. Panels B, D, F, and H show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%). The post-cutoff period (10/2023 onward) is shaded gray.
Figure 5: Recall of closing prices for AAPL, MSFT, and NVDA without context.
This figure shows the LLM’s estimated closing prices for AAPL, MSFT, and NVDA compared to the actual values. Panels A, C, and E graph the actual values against the estimated values. Panels B, D, and F show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%). The post-cutoff period (10/2023 onward) is shaded gray.
Figure 6: Recall of closing prices for AAPL, MSFT, and NVDA with context.
This figure shows the LLM’s estimated closing prices for AAPL, MSFT, and NVDA compared to the actual values. We give the LLM two previous end of the month closing prices given as context. Panels A, C, and E graph the actual values against the estimated values. Panels B, D, and F show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%). The post-cutoff period (10/2023 onward) is shaded gray.
Overall, these findings extend our market index results, highlighting that GPT-4o’s memorization is not uniform across securities and is sensitive to contextual cues and stock-specific factors. The high accuracy for META and GOOGL, especially with context, parallels the model’s strong recall of recent macroeconomic levels and index values, suggesting robust memorization of prominent, frequently referenced data. Conversely, larger errors for long-listed stocks like AAPL and MSFT, even with context, align with the weaker recall of long-horizon macroeconomic levels, pointing to the potential dilution of older data in the training corpus. This selective memorization reinforces the challenge for financial research: apparent forecasting success for individual stocks within the training period may reflect memorized prices rather than predictive ability, necessitating evaluations with post-cutoff data to ensure methodological rigor.
5.4.2 Individual Stocks in the Cross Section